OpenStudy (anonymous):

how to set this up On your next vacation, you will divide lodging between large resorts and small inns. Let x represent the number of nights spent in large resorts. Let y represent the number of nights spent in small inns. a) Write a system of inequalities that models the following conditions: you want to stay at least 5 nights. At least one night should be spent at a large resort. Large resorts average $200 per night and small inns average $100 per night. Your budget permits no more than $700 for lodging.

OpenStudy (anonymous):

Breakdown the question into smaller parts. The first inequality is the total number of nights must be equal to or greater than 5. How would you right that?

OpenStudy (anonymous):

*write

OpenStudy (anonymous):

200x+100<=700 which gives you 3 nights

OpenStudy (anonymous):

Hey Krissy, you are right about the first equation being 200x+100y<=700...I'm not entirely sure about the second one, and I'll need you to point out any flaws you might see in this, but I think it is something like 1<=x<x+y<=5

OpenStudy (anonymous):

okay and i'll tell you what i come up with

OpenStudy (anonymous):

okay i'm not sure how to get the last one at all

OpenStudy (anonymous):

is it x+y+>5, x=> 1, y=>0?

OpenStudy (anonymous):

i think that's right

OpenStudy (anonymous):

I think that x+y is greater than or equal to 5 and @Eyad I think I'm missing something :p

OpenStudy (anonymous):

oh i had the swithed around

OpenStudy (anonymous):

Sry i was afk :/ lemme read this question :D

OpenStudy (anonymous):

200x + 100y < = 700 at least one night in large resort (x) 1 >= x <= 5.....I am not sure where to put the y

OpenStudy (anonymous):

okay, could i also do this 3+y>=5 which makes y>=2

OpenStudy (anonymous):

wouldn't y be <= 4 ?

OpenStudy (anonymous):

5-3 is 2

OpenStudy (anonymous):

where are you getting the 3 from ?

OpenStudy (anonymous):

from the small resort when we got 3 as our anser for that and from there i just plug it into the x+y>=5 equation. not sure if i can do that though

OpenStudy (anonymous):

*answer

OpenStudy (anonymous):

This Answer will be x+y≥6 , x≥1,y≥0,250x+125y≤1400 .

OpenStudy (anonymous):

I wish if i can explain but Its very long besides i'am a bad teacher ,thats the answer just in case you want to relate .

OpenStudy (anonymous):

?

OpenStudy (anonymous):

your total is $700. $200 for large resort (x) $100 for small resort (y) stay at least 1 night for x.....1 >= x stay at the most 4 nights for y....y <= 4 200x + 100y = 700 I still dont see where you are getting the 3

OpenStudy (anonymous):

it says at least 5 nights not 4 i think that's where your messing up

OpenStudy (anonymous):

thanks anyways though guys for the help

OpenStudy (anonymous):

Oh...I thought it said ...staying 5 nights, not at least 5 nights. The reason I came up with 4 was because if she stays 5 nights, one (or possible more) would be at the small resort. So the small resort would be 4 or less. Sorry if I messed with your brain and confused you...my bad

OpenStudy (anonymous):

its okay